Solve each system of equations of substitution.
3x-y=8
x-y=8

Answer by ReadingBoosters(3246)

3x-y=8
x-y=8 or x=8+y
3(8+y)-y=8
24 + 3y - y = 8
24 + 2y = 8
2y = -16
y = -8
x = 8+y = 8+(-8)=0